# Programmable photonic waveguide arrays: opportunities and challenges

### v1

@@ Abstract
The field of programmable photonics has advanced significantly in recent decades, driven by the rising demand for complex applications, such as optical quantum computing and photonic neural networks. However, as the complexity of these applications increases, there is an increasing need for novel designs that enhance circuit transmission and enable further miniaturization. Photonic waveguide arrays (WAs) hold a unique position in integrated photonics, as they implement “always-on” Hamiltonians and have no direct analogs in free-space optics. They find applications in various fields, including light propagation studies, quantum walks, and topological photonics. Despite their versatility, the lack of reconfigurability has limited their utility and hindered further advancements for a long time. Recently, programmable waveguide arrays (PWAs) have emerged as a promising solution for overcoming the limitations of static WAs and PWA-based architectures have been proven to be universal. This perspective proposes a vision for photonic circuits based on PWAs as a new, interdisciplinary field. We review the history of the development of PWAs and outline their potential in areas such as simulation, communication, sensing, and classical and quantum information processing. This technology is expected to become increasingly feasible with advancements in programmable photonics, nanofabrication, and quantum control.
@@ I Introduction
The rapid advancements in photonics over the past few decades have been evident in numerous areas, from innovations in materials (such as silicon photonics [1, 2, 3, 4], III-V semiconductors [5, 6, 7]), to fabrication techniques (such as direct laser writing [8, 9] and thin film technologies [10, 11, 12]), and an expanding range of applications (including photonic quantum computing [13], topological photonics [14], quantum cryptography [15], and artificial intelligence [16]).
Programmable photonic circuits [17] enable the implementation of diverse functions within a single device, with applications spanning microwave photonics [18], neuromorphic computing [19], and quantum computing [20]. These devices primarily use mesh structures based on Mach-Zehnder interferometers (MZI) [21, 22, 23], relying on balanced (50:50) 2$\times$2 beam splitters and phase shifters, adapted from free-space optics. Devices employing these approaches are essential for emerging applications such as optical quantum computing and photonic neural networks. Fig. 1 summarizes the significant developments that have been seen over the past two decades. However, maintaining this rate of progress may be constrained by factors such as the need to bend the light paths to realize photonic circuits–contributing to photon loss and increasing the device footprint.
Waveguide arrays (WAs) play a key role in integrated photonics and have been utilized in various studies, including light propagation [24, 25, 26, 27], optical Bloch oscillations [28, 29], Anderson localization [30, 31], and wave dynamics in quasicrystals [32, 33]. They also have extensive applications in simulating physical phenomena [34, 14] and communication networks [35, 36]. In recent years, their applications have expanded further, particularly in the processing of classical and quantum information, including photonic matrix multiplication [37], quantum walks [38], quantum state manipulation [39, 40], and quantum state generation [41]. Fig. 2 summarizes the key demonstrations achieved with WA technology. WAs are formed from an array of evanescently coupled waveguides that confine the spatial area through which the light propagates. They have no direct analogs in free-space optics, enabling the implementation of “always-on” Hamiltonians [26] and providing direct access to higher-dimensional Hamiltonians. The development of these devices has a long history, as illustrated in Table. 2. However, until recently they have been typically static, single-purpose devices, lacking reconfigurability, and with their applications fixed upon fabrication.
Programmable waveguide arrays (PWAs) have recently emerged as a promising solution to address this challenge [42]. Such a device can realize programmable Hamiltonians and an architecture based on cascaded PWAs, such as the one illustrated in Fig. 3, offers a robust framework for implementing arbitrary unitary transformations [43]. Furthermore, it enables the simulation of both time-independent and time-dependent Hamiltonian evolution in a reconfigurable manner, which, combined with the potential of simulating higher-dimensional geometries [44], opens new applications in simulations and computing through an entirely different paradigm, serving as an attractive alternative to mesh structures [45, 46]. However, compared with mesh structures, PWA-based architectures do not require perfect fabrication for balanced beam splitters and can program light routes on-chip without bending the light paths [47]. This paves the way for a more compact photonic processor, with broad applications spanning photonic matrix multiplication [37] and topological photonics [48].
This perspective outlines a vision for photonic circuits based on PWAs, highlighting their potential impact (Table 1), addressing key challenges, and identifying future research directions. The structure begins with a review of applications that can be enhanced by PWAs in Sec. II, followed by an introduction to PWA device modeling and design in Sec. III. Device characterization and control are discussed in Sec. IV, and Sec. V concludes with final remarks and a future outlook.
Universal multiport interferometers [43]
1. Linear optical quantum computing [45, 20, 49] 2. Digital-analog quantum computation [46] 3. Photonic matrix multiplication [37] 4. Neuromorphic photonics (PNN) [50, 51, 52, 53, 54, 55, 56, 57, 58] 5. Quantum transport [59] 6. Photonic field-programmable gate arrays [60, 61] 7. Switching and data communication networks [62, 36] 8. Data encryption [63] 9. Microwave photonics [18]
Topological photonics [14, 48, 64, 65]
1. Optical communications [66] 2. Fault-tolerant quantum computing and spintronics [67] 3. Topological edge states generations [68, 69] 4. Generating robust quantum states [70, 71, 41, 39, 72, 73, 74, 75, 76] 5. Optical isolator [77, 78, 79] 6. Photonic Floquet topological insulators [80]
Nonlinear phenomena
1. Optical discrete solitons studies [81, 82, 24] 2. Optical discrete solitons in two-dimensional WAs [27] 3. Optical Bloch oscillations [28, 29] 4. Topological systems implementations [83, 84, 85, 86] 5. Optical isolation with topological photonics [87]
Other emerging applications
1. Quantum walk [88, 38, 89] 2. Simulating non-Markovian giant atom decay [90] 3. Simulating higher dimensional systems [44] 4. Anderson insulator [91, 92] 5. Implementing time-dependent Hamiltonian [80, 85, 93] 6. Non-Hermitian systems implementations [94, 95, 96, 97, 98] 7. Non-Abelian braiding [99, 100] 8. W-state generation [101, 102] 9. Perfect state transfer [103]
Table 1: Potential applications of PWA-based architectures. The references herein include demonstrations using static WAs and potential adaptations from the MZI-based discrete programmable scheme, as well as other applications that could be implemented with PWAs.
@@
Figure 1: Increasing complexity of programmable photonic circuits for advanced applications. The number of phase shifters is used for quantifying the tunable components. The development trend is indicated by an exponential function fitting represented by dashed lines. Notable milestones in the field of optical quantum computing (OQC) include the first static CNOT gate (a static device with 0 tunable components in 2008) [104], the first programmable quantum processor (8 tunable components in 2011) [105], the first universal linear circuits (30 tunable components in 2015) [20], the first large-scale circuit (93 tunable components in 2018) [106], and a recent large-scale circuit (216 tunable components in 2023) [107]. Key demonstrations in the field of photonic neural networks (PNN) include the first demonstration of unscrambling light (12 tunable components in 2017) [52], the first demonstration of deep learning (24 tunable components in 2017) [50], the first implementation of complex-valued neural networks (56 tunable components in 2021) [55], the first realization of backpropagation training for deep learning (30 tunable components in 2023) [57], the first demonstration of a fully-integrated coherent optical neural network (169 tunable components in 2024) [108].
Figure 2: Key demonstrations leveraging waveguide array technology. The observation of discrete spatial optical solitons [24], observation of linear and nonlinear optical Bloch oscillations [28], two-dimensional discrete solitons [27], demonstration of the evolution of linear and nonlinear waves in an implementation of the Anderson model [31], demonstration of quantum walk with correlated photons [38], observation of parity-time (PT) symmetry [95], demonstration adiabatic pumping via topological edge states [68], demonstration of a photonic Floquet topological insulator [80], demonstration of generating and verifying single-photon W-state [102], observation of a topological transition [96], demonstration of perfect state transfer of an entangled photonic qubit [103], demonstration of topologically bound states in a PT-symmetric dimerized waveguide array [97], demonstration of optical isolators based on breaking time reversal symmetry [79], demonstration of topological protection of biphoton states [41], demonstration of quantum interference of topological states of light [39], demonstration of topological light funneling [93], demonstration of Non-Abelian braiding [100], demonstration of the two-qubit optical controlled-NOT gate [40], demonstration programming Hamiltonian [42] and quantum logics [109]. The chronological order refers to the date of publication of the work.
Year Development Reference
1973 First experimental demonstration of a directional coupler fabricated in GaAs [110]
1997 First demonstration of a nonlinear waveguide array fabricated on an AlGaAs platform [111]
2002 Two-dimensional optically induced waveguide array in a bulk photorefractive crystal [25]
2013 Two-dimensional helical waveguides fabricated using femtosecond laser writing technology [80]
Recent Fully electro-optically programmable waveguide array in bulk lithium niobate [112, 42, 109]
Table 2: Evolution of waveguide array systems.
Figure 3: Schematics of an integrated PWA architecture. (a) The architecture incorporates photon sources and detectors with potential for integration as well as cascaded $n$ multisection PWAs. The black lines represent optical waveguides, and the orange lines indicate control connections. Each section of PWA has independent tuners and implements a $N$-dimensional unitary transformation matrix $U_{k}(\vec{v}_{k})$, where $\vec{v}_{k}$ is the control parameter applied to the tuners in the $k$-th section. The entire architecture implements an overall unitary transformation matrix $U$. For a given input state |$\ket{\psi_{\text{in}}}$, the output state is calculated as |$\ket{\psi_{\text{out}}}=U\ket{\psi_{\text{in}}}$. (Note: This architecture can be used for both classical and quantum light.) (b) Cross-section of an example PWA featuring three waveguides (black areas indicate the optical modes) based on the device recently demonstrated in Ref [42, 109]. The device is fabricated using annealed proton-exchange waveguides on x-cut lithium niobate [113]. Gold microelectrodes functioning as microtuners were patterned on top of the silicon dioxide buffer layer above the waveguides. Electric fields are generated by applying control pulses to the electrodes and are confined by the shielding effect of neighboring electrodes, effectively eliminating crosstalk [114]. This ensures precise control of individual Hamiltonian terms and their corresponding unitary transformations.
II Overview of potential applications
This section discusses applications of WA and MZI-based structures that may be transformed by adopting PWAs. Table. 1 provides a summary of the potential applications.
II.1 Universal multiport interferometers
Universal multiport interferometers enable arbitrary linear unitary matrix transformations between multiple optical channels. They have extensive applications in both classical and quantum information processing [17, 37], which will be discussed in more detail in the following content.
In addition to MZI-based mesh structures, which are considered discrete or digitized, there has been growing interest in leveraging continuously coupled waveguides or multimode interferometers (MMIs) as error-robust and more scalable alternatives to realize universal multiport interferometers [115, 116]. In 2017, a hybrid structure based on cascading MMIs and phase shifters was proposed to implement unitary transformations [117]. This structure was further investigated in [118], where several transformations were numerically studied. Experimental demonstrations of such an architecture have also been performed [119, 120]. Enhanced scalability and error tolerance have also been demonstrated for large-scale circuits [121]. The realization of multiport transformations using structures with partially programmable waveguide arrays has also been proposed [122]. However, these structures still require the decoupling and recoupling of waveguides or lack full programmability, which does not fully exploit the advantages of continuously coupled waveguides in integrated photonics.
Universality of PWAs
Structures based on cascaded PWAs take full advantage of continuously coupled waveguides in integrated photonics. They have also been recently rigorously proven to be universal [43], addressing a long-standing question raised in previous studies. The work introduced a decomposition algorithm for realizing any unitary matrix using PWAs with guaranteed error bound. For $U(2)$, a maximum of 4 sections are needed, which is the exact decomposition. The required number of sections $n$ scales as $O(N^{3}k)$, with the error scaling as $O(N^{3}/k)$, where $k$ is the Trotterization number. While this result suggests a high number of sections may be required to achieve low fidelities, the study demonstrates that, in practice, the number of sections needed is significantly smaller than the theoretical bound. By leveraging realistic materials parameters and numerical optimization to compile the unitary, these requirements become practically achievable. Therefore, cascaded PWAs are functionally equivalent to the widely adopted MZI-mesh structures [21, 22], while they eliminate the need for bending. This unique feature can potentially lead to significant reductions in optical losses and circuit footprint.
II.2 Quantum computing
Quantum computers promise to solve computational problems more efficiently than classical systems [123]. In 1982, Feynman proposed using controllable quantum systems to simulate the dynamics of other systems [124], now known as analog quantum computing [125, 126]. This approach allows the simulator’s Hamiltonian to be controlled to match the target system, offering greater error robustness, less demanding hardware, and better scalability compared to universal digital architectures [127, 128, 129]. Waveguide arrays, which naturally implement time evolution governed by their Hamiltonian, have been used to simulate optical analogs of solid-state systems [68]. By varying waveguide coupling along the propagation length and integrating control electrodes to tune coupling strength, PWAs can simulate time-dependent Hamiltonians for analog quantum simulations.
Integrated photonics is a promising platform for gate-based quantum circuits [130, 131]. Digital quantum computing [132], where qubits evolve through single- and two-qubit gates, is universal [133] and can be implemented in photonic systems using mesh structures of $2\times 2$ beam splitters and phase shifters [21, 22, 23]. While static, continuously coupled waveguides have been used to implement quantum gates [49, 40], digital quantum circuits require reconfigurable systems. Recently, controllable single-qubit gates have been demonstrated [109], and cascaded PWAs offer a pathway to universal digital quantum circuits [43]. A PWA-based digital quantum simulator could implement Trotterization steps by discretizing time evolution into small unitary operations [133]. By tuning the waveguide coupling and interaction parameters, this setup can approximate the continuous evolution of a system’s Hamiltonian, simulating digital quantum gates effectively. This approach provides a scalable route to implement digital quantum algorithms using cascaded waveguide arrays.
By supporting both analog and digital quantum information processing, cascaded PWAs offer a unique capability in photonics. Digital-analog quantum computing (DAQC), which combines the strengths of both approaches, has been proposed to address the challenges of the noisy intermediate-scale quantum era [46]. Using analog methods for continuous evolution and digital precision for error correction, DAQC enhances efficiency, error tolerance, and scalability, offering a promising path to solve complex quantum problems with fewer resources.
II.3 Photonic neural networks
Programmable waveguide arrays present a compelling alternative hardware platform for implementing photonic neural networks. Compared to MZI-based or hybrid MMI–phase shifter architectures, PWAs may more easily scale to larger input dimensions, support multi-wavelength operations, and reduce overall optical losses. Furthermore, due to the universality and robustness of PWAs, the training of PWA-based PNNs can be streamlined with simplified calibration process. Additionally, the continuously coupled waveguides in PWAs provide unique interfaces for integrating nonlinear activation elements.
II.4 Topological photonics
Over the past decade, topological photonics has emerged as a prominent field [14, 48, 64], inspired by the discovery of quantum Hall effects and topological insulators in condensed matter physics. WAs and their variants enable the direct simulation of tight-binding Hamiltonians that describe electron behavior in topological materials [65]. As a result, they have become popular platforms for demonstrating intriguing phenomena, including the first experimental observation of PT symmetry [95].
Robustness is one of the many valuable aspects of topological photonics [136]. As recently demonstrated, PWAs can showcase multiple topological effects within a single device and evaluate system robustness by introducing noise through their reconfigurability [42]. Tunable optical isolators based on topological effects can also be achieved [78, 79]. Additionally, their application can be extended to the modeling of higher-dimensional systems using a mapping scheme [44] such as simulating non-Markovian giant atom decay [90]. Other promising applications include optical communications [66], spintronics [67] and quantum computing and with a wide range of demonstrations in generating topologically protected quantum states [70, 71, 41, 39, 72, 73, 74, 75, 76, 137].
Among the recent advancements in the field, photonic Floquet topological insulators are particularly prominent [80]. These two-dimensional WAs exhibit unique properties by utilizing helical waveguides to mimic the rotation of electrons in time as light propagates through them. By adding reconfigurability to induce disorder, their functionalities can be expanded to tunable Anderson insulators [91, 92]. Moreover, similar structures may be achieved by cascading multiple two-dimensional PWAs with dynamic modulation that allows the implementation of time-dependent Hamiltonians [93, 85] to mitigate the propagation loss in the helical design caused by the bending of the waveguides.
II.5 Nonlinear phenomena
Nonlinear phenomena extend the capabilities of photonic systems, exhibiting particularly intriguing features in coupled waveguides. The study of nonlinearity began with directional couplers [81] and later expanded to WAs [82]. Subsequent research demonstrated the effects of nonlinear phenomena on solitons [111, 24] and optical Bloch oscillations [28, 29], including their behavior in two-dimensional WAs [25, 27]. Leveraging the combined effects of nonlinearity and reconfigurability in PWAs is compelling, as these programmable structures offer a dynamic approach to precisely tune nonlinear effects for advanced light manipulation. Beyond providing flexibility in the nonlinear activation functions for PNNs, this synergy can greatly enhance applications such as dynamic pulse shaping [138] and all-optical switching and routing [139].
With the growing interest in topological photonics, research on nonlinearity has expanded to encompass advanced topological systems utilizing WAs [83, 84, 85, 86]. The integration of nonlinear effects with topological phenomena unlocks new possibilities, including topological lasers that are robust against fabrication imperfections [87], as well as the on-chip generation of topologically protected entangled photon pairs via nonlinear processes like spontaneous parametric down-conversion (SPDC) and four-wave mixing [140, 141, 142, 143]. Incorporating reconfigurability can further enhance these applications. PWAs leveraging materials with nonlinear properties can enable precise light filtering and improve wavelength conversion efficiency.
II.6 Other emerging applications
The incorporation of gain and loss in waveguide arrays enables the realization and control of non-Hermitian dynamics [94], unlocking phenomena such as parity-time symmetry [95] and topological phase transitions [96], as well as applications like the generation of topologically protected bound states [97]. In PWA systems, these effects can be achieved by applying various control mechanisms to waveguide tuners, including the creation of defect waveguides [144, 98]. Additionally, non-Abelian braiding, a key concept in topological physics, has recently been proposed and demonstrated in photonic WA systems [99, 100]. PWAs could offer a dynamic platform for emulating and harnessing non-Abelian braiding, paving the way for topological quantum computing [145, 64].
Additional potential applications of precisely engineered device Hamiltonians and lengths include quantum walks [88, 38, 89], simulating non-Markovian giant atom decay [90] and higher-dimensional systems [44], generating W states [101, 102], and achieving perfect state transfer [103]. The reconfigurability of PWAs offers a unified platform for realizing these applications with enhanced precision, scalability, and adaptability. Time-dependent Hamiltonians could potentially be realized by cascading PWAs, with each section implementing a different Hamiltonian [80, 85, 93].
III Device modeling and design
@@
In this section, we explore the challenges in PWA modeling and design, along with potential strategies to address them. We begin with a brief overview of waveguide array systems, followed by two mathematical modeling approaches for PWA devices. Lastly, we review photonic material platforms suitable for realizing PWAs and address essential considerations for circuit layout and waveguide geometry design.
III.1 Overview of continuously coupled waveguide systems
Traditional WAs are constructed by using arrays of evanescently coupled waveguides on photonic platforms. These devices allow optical waves to spread and propagate through. Owing to the simplicity of their fabrication, they were initially proposed to achieve light discretization [146]. Since the first theoretical proposal in 1965 [147] and experimental demonstrations in 1973 [110], WAs have been fabricated on various material platforms along with advancements in their device size and circuit complexity. Table. 2 highlights the key milestones in their evolution. The smallest architecture of PWA is a reconfigurable directional coupler corresponding to two coupled waveguides.

### v3

@@ Abstract
The rising complexity of photonic applications, ranging from quantum computing to neuromorphic processing, has driven the demand for highly programmable and scalable photonic integrated circuits. While mesh-based architectures built from Mach-Zehnder interferometers have enabled significant advances, their reliance on beam splitting and light bending introduces optical loss, fabrication challenges, and scalability bottlenecks. Continuously lateral-coupled integrated waveguide arrays (WAs), by contrast, offer compact systems with no direct free-space analogs, but their static nature has limited their utility. Recently, programmable waveguide arrays (PWAs) have emerged as a promising alternative, combining the Hamiltonian richness of WAs with tunable control. This perspective outlines the conceptual foundations, recent progress, and future potential of PWAs across quantum simulation, photonic neural networks, topological photonics, and nonlinear optics. We examine the theoretical and practical challenges of modeling, fabrication, and control, and propose PWAs as a next-generation architecture for compact, reconfigurable photonic processors.
@@ I Introduction
The rapid progress in photonic technologies over the past few decades has transformed both research and industry, driven by advances in materials (e.g., silicon photonics [1, 2, 3, 4], III-V semiconductors [5, 6, 7]), fabrication techniques (e.g., direct laser writing [8, 9], thin-film technologies [10, 11, 12]), and applications ranging from quantum computing [13] to topological photonics [14], quantum cryptography [15], and artificial intelligence [16].
Among these advances, programmable photonic circuits have become a key enabler of flexible and scalable devices [17], with Mach-Zehnder interferometer (MZI) based mesh structures [18, 19, 20] being widely adopted for applications in microwave photonics [21], neuromorphic computing [22], and quantum information processing [23]. These mesh architectures rely on discrete 2×2 beam splitters and phase shifters, adapted from free-space optics, and have scaled rapidly in complexity over the last two decades, as illustrated in Fig. 1, which summarizes the growth in circuit size and component count across different photonic platforms. However, they are fundamentally limited by the need to bend light paths, contributing to optical loss and limiting circuit density.
In contrast, photonic waveguide arrays (WAs) offer a unique platform that implements “always-on” Hamiltonians through evanescent coupling in a fixed spatial geometry. These structures have no direct analogs in free-space optics and have enabled pioneering studies in light transport [24, 25, 26, 27], Bloch oscillations [28, 29], Anderson localization [30, 31], quasicrystal dynamics [32, 33], and topological photonics [14, 34]. Their use has also expanded to classical and quantum information processing, including quantum walks [35], matrix multiplication [36], and quantum state generation [37, 38, 39]. Traditional waveguide arrays have long enabled studies in light dynamics and quantum simulations. However, their static nature, with parameters fixed at fabrication, imposes severe limitations on scalability, adaptability, and reusability across applications.
We propose a new paradigm: programmable waveguide arrays (PWAs) [40], which merge the continuous coupling of static arrays with the reconfigurability of mesh networks [41]. This hybrid architecture opens unprecedented avenues in quantum computing, neuromorphic photonics, and topological protection [42, 43]. Figure 2 illustrates the conceptual architecture of programmable waveguide arrays, highlighting how local control elements enable dynamic reconfiguration of coupling and on-site parameters. Beyond active reconfiguration, PWAs also leverage structural programmability: the coupling between adjacent guides can be preset during fabrication by controlling their separation, enabling fine-grained tuning of the device’s Hamiltonian even before dynamic control is applied. Bridging concept and implementation, recently developed PWAs introduce tunability through these local controls [40], allowing direct programming of Hamiltonians and the realization of arbitrary unitary transformations without resorting to light-bending geometries. This shift marks a departure from conventional MZI-based processors, offering a compact and intrinsically analog alternative that could scale more favourably for large or dynamically evolving systems. Furthermore, PWA-based architectures have recently been shown to be universal [44], capable of emulating both time-independent and time-dependent Hamiltonians with reduced optical loss and minimal footprint.
In this perspective, we highlight the emerging potential of continuously-coupled integrated programmable waveguide arrays as a versatile interdisciplinary platform. We start with an introduction to their operating principles and examine the associated challenges in modeling and design (Sec. II). This is followed by a discussion on device characterization and control (Sec. III). We then survey application domains where PWAs could unlock new capabilities (Sec. IV. Finally, we conclude with a future outlook, presenting a roadmap and identifying open research directions (Sec. V).
@@
Figure 1: Increasing complexity of programmable photonic circuits for advanced applications, quantified by the number of tunable phase shifters per chip. The development trend is indicated by an exponential function fitting represented by dashed lines. Although recent milestones have been achieved primarily with MZI mesh networks, the continued exponential growth in complexity highlights the demand for alternative architectures. PWAs may provide a more scalable and compact solution to continue this trend. Notable milestones in the field of optical quantum computing (OQC) include the first static CNOT gate (a static device with 0 tunable components in 2008) [45], the first programmable quantum processor (8 tunable components in 2011) [46], the first universal linear circuits (30 tunable components in 2015) [23], the first large-scale circuit (93 tunable components in 2018) [47], and a recent large-scale circuit (216 tunable components in 2023) [48]. Key demonstrations in the field of photonic neural networks (PNN) include the first demonstration of unscrambling light (12 tunable components in 2017) [49], the first demonstration of deep learning (24 tunable components in 2017) [50], the first implementation of complex-valued neural networks (56 tunable components in 2021) [51], the first realization of backpropagation training for deep learning (30 tunable components in 2023) [52], the first demonstration of a fully-integrated coherent optical neural network (169 tunable components in 2024) [53].
[…]
Figure 3: Evolution of waveguide array systems. (a) The first experimental demonstration of a directional coupler fabricated in GaAs in 1973 [56]. (b) The first demonstration of a nonlinear waveguide array fabricated on an AlGaAs platform in 1997 [57]. (c) The two-dimensional optically induced waveguide array in a bulk photorefractive crystal in 2002 [25]. (d) The two-dimensional helical waveguides fabricated using femtosecond laser writing technology in 2013 [58]. (e) The fully electro-optically programmable waveguide array fabricated in bulk lithium niobate used in recent demonstrations [59, 40, 41] (Red light shines through the chip to highlight the device. Credit to photographer: James Thomas).
II Device modeling and design
@@
In this section, we introduce the fundamentals of PWA and examine the challenges in PWA modeling and design, along with potential strategies to address them. We begin with a brief overview of waveguide array systems, followed by two mathematical modeling approaches for PWA devices. Lastly, we review photonic material platforms suitable for realizing PWAs and address essential considerations for circuit layout and waveguide geometry design.
II.1 Overview of continuously coupled waveguide systems
Traditional waveguide arrays are composed of evanescently coupled waveguides integrated on photonic platforms. These devices enable optical waves to discretize and propagate through coupled channels, a concept first proposed to achieve light discretization [60]. Since the initial theoretical proposal in 1965 [61] and the first experimental demonstrations in 1973 [56], WAs have been realized on various material platforms, with steady advances in device size and circuit complexity. The smallest PWA architecture is a reconfigurable directional coupler–two coupled waveguides augmented by a local control element that enables dynamic tuning of their coupling. The field is now advancing toward more complex configurations involving multi-waveguide and multi-section PWAs. The conceptual evolution of waveguide arrays is illustrated in Fig. 3, tracing the progression from rigid, fabrication-limited lattices to dynamically tunable architectures that define programmable waveguide arrays. This transition marks a shift from passive photonic design to active, software-defined photonic functionality.